Calculate the following properties for the given object. Show your work:

a) Specific Weight

b) Specific Gravity

4 cm Mass 72 g 11 cm
Answer #1

1) specific weight is weeight for unit volume of material = weight/volume=mass*g / volume

Volume= area*height=πr^2*h=3.14*4*11=138.16 cm^3

Density=mass/volume=72 /138.16=0.5211 gram/cc

Specific weight=Density*g=0.522*10^3 kg/m^3 * 9.8 =5107.122 kg/(m^2s^2)

2)specific gravity =density of substance/density of water= 0.5211/1=0.5211